ROTATIONAL SPECTRA AND STRUCTURES OF (OCS)_{2}CO_{2} and (OCS)_{3}

\underline{REBECCA A. PEEBLES}, SEAN A. PEEBLES and ROBERT L. KUCZKOWSKI, University of Michigan, Department of Chemistry, 930 N. University Ave., Ann Arbor, MI 48109.

The structures of (OCS)_{2}CO_{2} and (OCS)_{3} have been determined from rotational spectra obtained using a pulsed nozzle Fourier transform spectrometer. The two trimers have a barrel like monomer arrangement commonly seen in similar complexes. In (OCS)_{2}CO_{2}, the two OCS units have an anti-parallel arrangement, like the OCS dimer, but twisted from planarity by 34^{∘}. The CO_{2} straddles both of them in an approximately parallel orientation. In (OCS)_{3}, two of the OCS molecules are again nearly anti-parallel with the third OCS straddling them. Seven isotopic species were assigned for (OCS)_{2}CO_{2}. Two isotopic species were assigned for (OCS)_{3}, to supplement the normal species data previously reported.^{a} The dipole moments of the trimers are consistent with the derived structures. Comparisons of the structures with results from a semi-empirical model including electrostatic, dispersion and repulsive terms will be presented.

aJ.P. Connelly, A. Bauder, A. Chisholm and B. J. Howard, Mol. Phys. 88 (1996) 915